OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

I already factored it into \[7(sinx)^2=0\] i dont know what to do now

OpenStudy (anonymous):

Is the answer \[\pi/2\]?

OpenStudy (mathmate):

The factorization is not exactly correct. You can substitute \(s=sin(x)\) to give \[7s^2-14s+7=0\] which factorizes to \[s^2-2s+1=0\] Can you solve for s?

OpenStudy (mathmate):

or even better: \[(s-1)^2=0\] can you solve for s, and hence x?

OpenStudy (anonymous):

(x-1)^2

OpenStudy (anonymous):

oh wait i meant (sinx-1)^2

OpenStudy (mathmate):

That's great! Continue!

OpenStudy (anonymous):

i dont know what to do now

OpenStudy (mathmate):

Hint: